OpenStudy (anonymous):

what's the graphic of y = (-1.2)^x

OpenStudy (anonymous):

just a sketch of what it will look like

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

@robtobey, but when x is -1.5, isn't (-1.2)^(-1.5) impossible? also, in y = a^x, isn't 'a' should be at least 0 ? math geniuses, help plzz..

OpenStudy (anonymous):

(-1.2)^(-1.5) is imaginary so if you don't know complex numbers then you can't do it. In y = a^x, why should a be at least 0?

OpenStudy (anonymous):

@remnant well simply because if 'a' is negative, for example -1.2. When x is 1.5, y = (-1.2)^(1.5), which as you said, is imaginary

OpenStudy (anonymous):

however, if you look at the graph on Wolfram Alpha, http://www.wolframalpha.com/input/?i=y+%3D+%28-1.2%29%5E%28x%29 the negative part also has blue lines, which they say is 'real part'.. im so confused :S

OpenStudy (anonymous):